### v1

@@ Appendix B Quadratic program and linear program
$({\rm QP.D})\quad\begin{array}[]{|ll}\underset{\bm{\theta}_{1}\in\mathbb{R}^{n},\bm{\theta}_{2}\in\mathbb{R}^{p}}{\mbox{maximize}}&\bm{b}^{\top}\bm{\theta}_{1}-\frac{1}{2}\bm{\theta}_{2}^{\top}\bm{Q}\bm{\theta}_{2}\\ \mbox{subject to}&\bm{A}^{\top}\bm{\theta}_{1}-\bm{Q}\bm{\theta}_{2}\leq\bm{c},\\ \end{array}$
@@
$({\rm QP.D}^{\prime})\quad\begin{array}[]{|ll}\underset{\bm{\theta}_{1}\in\mathbb{R}^{n},\bm{\theta}_{2}}{\mbox{minimize}}&f(\bm{\theta}_{1},\bm{\theta}_{2})=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1})+\frac{1}{2}\bm{\theta}_{2}^{\top}\hat{\bm{Q}}\bm{\theta}_{2}\\ \mbox{subject to}&\bm{c}-\bm{A}^{\top}\bm{\theta}_{1}+\bm{Q}\bm{\theta}_{2}\geq\bm{0}.\end{array}$
We see that (QP.D) is equivalent to (QP.D) since $f(\bm{\theta}_{1},\bm{\theta}_{2})=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1})+\frac{1}{2}\bm{\theta}_{2}^{\top}\hat{\bm{Q}}\bm{\theta}_{2}=-\frac{1}{S}\left(\bm{b}^{\top}\bm{\theta}_{1}-\frac{1}{2}\bm{\theta}_{2}^{\top}\bm{Q}\bm{\theta}_{2}\right)+\frac{1}{S}\bm{\xi}^{\top}\bm{c}=-\frac{1}{S}``$objective of (QP.D) ”$+$“constant.”
@@
$p_{\bm{x}}(\bm{\theta}_{1},\bm{\theta}_{2})$ $=p_{\bm{x}_{1}}(\bm{\lambda})\,p_{\bm{x}_{2}}(\bm{\theta}_{2})=p_{\bm{x}_{1}}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1}+\bm{Q}\bm{\theta}_{2})\,p_{\bm{x}_{2}}(\bm{\theta}_{2})=\hat{p}_{\bm{x}_{1}}(\bm{\theta}_{1},\bm{\theta}_{2})\,p_{\bm{x}_{2}}(\bm{\theta}_{2}),$ (82)
@@
It is easy to see that (82) is a member of the exponential family.
@@
the condition (54) for (QP.D) is fulfilled with $p_{\bm{x}}(\bm{\theta}_{1},\bm{\theta}_{2})$ defined by (82), $f(\bm{\theta}_{1},\bm{\theta}_{2})=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1})+\frac{1}{2}\bm{\theta}_{2}^{\top}\hat{\bm{Q}}\bm{\theta}_{2}$, and $\Theta=\mathbb{R}^{p}_{+}$.
@@
[…]
$=\exp\big{(}-\hat{\bm{\xi}}^{\top}\bm{\lambda}\big{)}.$ (83)
@@
For $(\bm{\theta},\bar{\bm{\theta}})\in\mathrm{int}\Theta\times\mathrm{int}\Theta$, the surrogate $u(\bm{\theta}|\bar{\bm{\theta}})$ is a convex function of the form:
@@
$(\bm{\theta}_{1}^{(t)},\bm{\theta}_{2}^{(t)})\leftarrow\arg\min_{(\bm{\theta}_{1},\bm{\theta}_{2})}\left\{\begin{array}[]{r}-\sum\limits_{j\in[p]}\Big{(}\bm{a}_{j}^{\top}\bm{\theta}_{1}-\bm{q}_{j}^{\top}\bm{\theta}_{2}+\mu^{(t-1)}_{j}\ln(c_{j}-\bm{a}_{j}^{\top}\bm{\theta}_{1}+\bm{q}_{j}^{\top}\bm{\theta}_{2}\Big{)}\\ +\frac{1}{2}\bm{\theta}_{2}^{\top}\bm{\Sigma}^{-1}\bm{\theta}_{2}+(\bm{\nu}^{(t-1)})^{\top}\bm{\theta}_{2}\end{array}\right\},$ (86)
@@
$({\rm LP.D}^{\prime})\quad\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{n}}{\mbox{minimize}}&f(\bm{\theta}):=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1})\\ \mbox{subject to}&\bm{c}-\bm{A}^{\top}\bm{\theta}\geq\bm{0}.\end{array}$
@@
For $p_{\bm{x}}(\bm{\lambda})=\hat{p}_{\bm{x}}(\bm{\theta}):=\prod_{j\in[p]}\frac{\big{(}c_{j}-\bm{a}_{j}^{\top}\bm{\theta}\big{)}^{x_{j}}\exp\big{(}-(c_{j}-\bm{a}_{j}^{\top}\bm{\theta}_{1})\big{)}}{x_{j}!}$ and $G$ is defined by $G(\bm{x}):=\exp\Big{(}\sum_{j\in[p]}x_{j}\ln(1-\hat{\xi}_{j})\Big{)}$, the condition (54) for (LP.D) is fulfilled with $p_{\bm{x}}(\bm{c}-\bm{A}^{\top}\bm{\theta})$ and $f(\bm{\theta})=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1})$ and $\Theta=\mathbb{R}^{p}_{+}$.
@@ Appendix C EM gradient algorithms for the extended formulations
Algorithm 8 shows an EM gradient algorithm for the subproblem (16).
Algorithm 8 An EM gradient algorithm for (53)
1:Let $\bm{\theta}^{(0)}\in\mathrm{int}\Theta$, $\bar{f}\leftarrow f(\bm{\theta}^{(0)})(\equiv u(\bm{\theta}^{(0)}|\bm{\theta}^{(0)}))$ and $\beta\in(0,1)$.
@@
$\hat{\bm{\theta}}\leftarrow\bm{\theta}^{(t-1)}-\alpha\left(\nabla^{2}u(\bm{\theta}|\bm{\theta}^{(t-1)})\big{|}_{\bm{\theta}=\bm{\theta}^{(t-1)}}\right)^{-1}\left.\nabla u(\bm{\theta}|\bm{\theta}^{(t-1)})\right|_{\bm{\theta}=\bm{\theta}^{(t-1)}}.$ (87)
@@
5:         $\alpha\leftarrow\alpha\beta$ and update $\hat{\bm{\theta}}$ via (87), and set $\bar{f}\leftarrow f(\hat{\bm{\theta}})$ if $\hat{\bm{\theta}}\in\mathrm{int}\Theta$ and $f(\hat{\bm{\theta}})<\bar{f}$.
@@
In the plain vanilla EM gradient, the Newton update can go out of the domain of $u$, so a backtracking procedure is employed for the inner loop (lines 4 to 6) to ensure the update to be strictly feasible to the original problem and the objective value to decrease. Note that due to the differentiability of $u$ and $f$ and the first-order surrogate condition, this inner loop will end in a finite number of backtracking.

### v2

@@ Appendix B Quadratic program and linear program
[…]
@@
$({\rm QP.D}^{\prime})\quad\begin{array}[]{|ll}\underset{\bm{\theta}_{1}\in\mathbb{R}^{n},\bm{\theta}_{2}\in\mathbb{R}^{p}}{\mbox{minimize}}\quad&f(\bm{\theta}_{1},\bm{\theta}_{2})=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1})+\frac{1}{2}\bm{\theta}_{2}^{\top}\hat{\bm{Q}}\bm{\theta}_{2}\\ \mbox{subject to}&\bm{c}-\bm{A}^{\top}\bm{\theta}_{1}+\bm{Q}\bm{\theta}_{2}\geq\bm{0}.\end{array}$
We see that (QP.D) is equivalent to (QP.D) since $f(\bm{\theta}_{1},\bm{\theta}_{2})=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1})+\frac{1}{2}\bm{\theta}_{2}^{\top}\hat{\bm{Q}}\bm{\theta}_{2}=-\frac{1}{S}\left(\bm{b}^{\top}\bm{\theta}_{1}-\frac{1}{2}\bm{\theta}_{2}^{\top}\bm{Q}\bm{\theta}_{2}\right)+\frac{1}{S}\bm{\xi}^{\top}\bm{c}=-\frac{1}{S}$``objective of (QP.D) ”++“constant.”
@@
$p_{\bm{x}}(\bm{\theta}_{1},\bm{\theta}_{2})$ $=p_{\bm{x}_{1}}(\bm{\lambda})\,p_{\bm{x}_{2}}(\bm{\theta}_{2})=p_{\bm{x}_{1}}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1}+\bm{Q}\bm{\theta}_{2})\,p_{\bm{x}_{2}}(\bm{\theta}_{2})=\hat{p}_{\bm{x}_{1}}(\bm{\theta}_{1},\bm{\theta}_{2})\,p_{\bm{x}_{2}}(\bm{\theta}_{2}),$ (B.1)
@@
It is easy to see that (B.1) is a member of the exponential family.
@@
the condition (5.6) for (QP.D) is fulfilled with $p_{\bm{x}}(\bm{\theta}_{1},\bm{\theta}_{2})$ defined by (B.1), $f(\bm{\theta}_{1},\bm{\theta}_{2})=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta}_{1})+\frac{1}{2}\bm{\theta}_{2}^{\top}\hat{\bm{Q}}\bm{\theta}_{2}$, and $\Theta=\mathbb{R}^{p}_{+}$.
@@
$=\prod_{j\in[p]}\exp\big(-\hat{\xi}_{j}\lambda_{j}\big)$ $(\because\mbox{ MGF with }t=\ln(1-\hat{\xi}_{j}))$
$=\exp\big(-\hat{\bm{\xi}}^{\top}\bm{\lambda}\big).$ (B.2)
@@
For $\bm{\theta},\bar{\bm{\theta}}\in\mathrm{int}\{\bm{\theta}=(\bm{\theta}_{1},\bm{\theta}_{2})|\bm{A}^{\top}\bm{\theta}_{1}-\bm{Q}\bm{\theta}_{2}\leq\bm{c}\}$, the surrogate $u(\bm{\theta}|\bar{\bm{\theta}})$ is a convex function of the form:
@@
$(\bm{\theta}_{1}^{(t)},\bm{\theta}_{2}^{(t)})\leftarrow\arg\min_{(\bm{\theta}_{1},\bm{\theta}_{2})}\left\{\begin{array}[]{r}-\sum\limits_{j\in[p]}\Big(\bm{a}_{j}^{\top}\bm{\theta}_{1}-\bm{q}_{j}^{\top}\bm{\theta}_{2}+\mu^{(t-1)}_{j}\ln(c_{j}-\bm{a}_{j}^{\top}\bm{\theta}_{1}+\bm{q}_{j}^{\top}\bm{\theta}_{2}\Big)\\ +\frac{1}{2}\bm{\theta}_{2}^{\top}\bm{\Sigma}^{-1}\bm{\theta}_{2}+(\bm{\nu}^{(t-1)})^{\top}\bm{\theta}_{2}\end{array}\right\},$ (B.5)
@@
$({\rm LP.D}^{\prime})\quad\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{n}}{\mbox{minimize}}\quad&f(\bm{\theta}):=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta})\\ \mbox{subject to}&\bm{c}-\bm{A}^{\top}\bm{\theta}\geq\bm{0}.\end{array}$
@@
For $p_{\bm{x}}(\bm{\lambda})=\hat{p}_{\bm{x}}(\bm{\theta}):=\prod_{j\in[p]}\frac{\big(c_{j}-\bm{a}_{j}^{\top}\bm{\theta}\big)^{x_{j}}\exp\big(-(c_{j}-\bm{a}_{j}^{\top}\bm{\theta})\big)}{x_{j}!}$ and $G$ is defined by $G(\bm{x}):=\exp\Big(\sum_{j\in[p]}x_{j}\ln(1-\hat{\xi}_{j})\Big)$, the condition (5.6) for (LP.D) is fulfilled with $p_{\bm{x}}(\bm{c}-\bm{A}^{\top}\bm{\theta})$ and $f(\bm{\theta})=\hat{\bm{\xi}}^{\top}(\bm{c}-\bm{A}^{\top}\bm{\theta})$ and $\Theta=\mathbb{R}^{p}_{+}$.
@@ Appendix C EM gradient algorithms for the extended formulations
Algorithm 8 shows an EM gradient algorithm for the subproblem (3.6).
Algorithm 8 An EM gradient algorithm for (5.5)
1:Let $\bm{\theta}^{(0)}$ be an interior point of the feasible region of (5.5), $\bar{f}\leftarrow f(\bm{\theta}^{(0)})(=u(\bm{\theta}^{(0)}|\bm{\theta}^{(0)}))$ and $\beta\in(0,1)$.
@@
$\hat{\bm{\theta}}\leftarrow\bm{\theta}^{(t-1)}-\alpha\left(\nabla^{2}u(\bm{\theta}|\bm{\theta}^{(t-1)})\big|_{\bm{\theta}=\bm{\theta}^{(t-1)}}\right)^{-1}\left.\nabla u(\bm{\theta}|\bm{\theta}^{(t-1)})\right|_{\bm{\theta}=\bm{\theta}^{(t-1)}}.$ (C.1)
@@
[…]
@@
In the vanilla EM gradient, the Newton update can go out of the domain of $u$, so a backtracking procedure is employed for the inner loop (lines 4 to 6) to ensure the update to be strictly feasible to the original problem and the objective value to decrease. Note that due to the differentiability of $u$ and $f$ and the first-order surrogate condition, this inner loop will end in a finite number of backtracking.